OpenStudy (anonymous):

Write the slope equation for the line that passes through (-1,-4) and (4,11). Show all work please

OpenStudy (anonymous):

I was practicing a very fast method for this in the evening so am going to give this answer straight away with a probability of being wrong :P so here is it 3y=x+11

OpenStudy (karatechopper):

y=14/5x+b y+4=14/5(x+1) 5(y+4=14/5(x+1)) 5y+20=14(x+1) 5y+20=14x+14 5y+20-20=14x+14-20 5y=14x-6 5y/5=14x-6/5 y=14/5x-6/5

OpenStudy (karatechopper):

not correct FFM srry

OpenStudy (anonymous):

Yes, it is actually $$-y+3x=1$$